OpenStudy (anonymous):

Factor 16x2 + 24x + 9.

OpenStudy (callisto):

Use the identity (a+b)^2 = a^2 + 2ab + b^2 16x^2 + 24x + 9 = (4x)^2 + 2(4x)(3) + 3^2 = ... ?

OpenStudy (anonymous):

answers (16x + 3)(x + 3) (4x + 3)(4x - 3) (4x + 3)^2

OpenStudy (callisto):

Can you try first?

OpenStudy (anonymous):

the second one

OpenStudy (callisto):

Nope.... As I've said, use the identity (a+b)^2 = a^2 + 2ab + b^2 From 16x^2 + 24x + 9 = (4x)^2 + 2(4x)(3) + 3^2, can you deduce what a and b is

OpenStudy (anonymous):

what does deduce mean

OpenStudy (callisto):

deduce => guess

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4 and 3

OpenStudy (anonymous):

16x2 + 24x +9 16x^2 and 9 are both perfect squares...0f 4x and 3 twice their product is 24 x ,so the polynomial fits condition of being a perfect sq. 16x^2 + 24x 9 = (4x + 3)^2

OpenStudy (callisto):

3 is correct for b 4 is not correct though ...

OpenStudy (anonymous):

so t has to be 16

OpenStudy (anonymous):

3 and 16

OpenStudy (callisto):

Compare the general form and the things you get \[a^2 + 2ab + b^2\]\[ (4x)^2 + 2(4x)(3) + 3^2\] can you see what is a and b clearly? You've made a right guess for b... what is a?

OpenStudy (anonymous):

2

OpenStudy (callisto):

Nope..

OpenStudy (anonymous):

then it has to be 4

OpenStudy (callisto):

It has to be 4x :P

OpenStudy (anonymous):

oh ok

OpenStudy (callisto):

Put a= 4x and b = 3 into (a+b)^2. You can get your answer~

OpenStudy (anonymous):

ok thank you